OpenStudy (anonymous):

sqrt{x+5} = sqrt{2x - 1} solve?

OpenStudy (anonymous):

just kidding, copied and pasted wrong thing sorry!

OpenStudy (anonymous):

square both saide (dqrt(x+5))^2 = (sqrt(2x-1))^2 x+5=2x-1 x+5-x=2x-1-x 5=x-1 5+1=x-1+1 x=6

OpenStudy (anonymous):

\[\sqrt{5x-9}-3=\sqrt{x+4}-2\]

OpenStudy (anonymous):

^right one

OpenStudy (anonymous):

rigth one is a pain

OpenStudy (anonymous):

\[\sqrt{5x-9}-3=\sqrt{x+4}-2\] \[\sqrt{5x-9}=\sqrt{x+4}+1\] now square both sides

OpenStudy (anonymous):

\[5x-9=(\sqrt{x+4}+1)^2=x+4+1+2\sqrt{x+4}=x+5+\sqrt{x+4}\]

OpenStudy (anonymous):

subtract x and 5 from both sides, get \[4x-14=\sqrt{x+4}\] and then square again!!

OpenStudy (anonymous):

sqrt(5x-9)-3 = sqrt(x+4)-2 sqrt(5x-9)-3+3 = sqrt(x+4)-2+3 sqrt(5x-9)=sqrt(x+4)+1 square both sides 5x-9=x+4 +2sqrt(x+4) +1 4x-14 = 2 sqrt (x+4) 2x-7=sqrt(x+4) square both sides (2x-7)^2= x+4 4x^2 -28x+49 = x+4 4x^2 -29x + 45=0 (x-5)(4x-9)=0 x=5 or x=9/4

OpenStudy (anonymous):

CoCoTsoi has it. i forgot the 2 in front of the \(\sqrt{x+4}\)

OpenStudy (anonymous):

satellite, you are good too. :D

OpenStudy (anonymous):

don't forget to check your answers, because when you square you could introduce an extraneous solution

OpenStudy (anonymous):

what's an extraneous solution?

OpenStudy (anonymous):

and I don't get your 3rd post, satellite73